| TIP: Click on subject to list as thread! | ANSI |
| echo: | |
|---|---|
| to: | |
| from: | |
| date: | |
| subject: | 5\02 Pt 1 ISS On-Orbit Status 02-05-2003 |
This Echo is READ ONLY ! NO Un-Authorized Messages Please!
~~~~~~~~~~~~~~~~~~~~~~~~ ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
03 May 2003
ISS On-Orbit Status 2 May 2003
Part 1 of 4
Day 160 for Expedition 6 (since STS-113 launch on 11/23/02 [158 days
aboard ISS]). Last-but-one day for Expedition 6 on orbit.
All ISS systems continue to function nominally, except as noted
previously or below.
Emphasis shifted today from handovers between the crews to
preparations for return and stowage of equipment on Soyuz 5S,
scripted carefully by TsUP beforehand in a detailed list of return
hardware and its exact location in the Soyuz-211/TMA-1 Descent
Module (DM).
MCC-M/TsUP reported this morning that packing has gone well and
should be finished by end of the working day. The crew has
successfully completed two descent training sessions, wrapped up
"even with a signed protocol statement" by the Descent Commission.
CDR-7 Yuri Malenchenko continued the Russian BTKh-10 "Conjugation"
experiment, which pursues development of new methods of engineering
of new recombinant strains of protein producers. [Yuri's job again
was to monitor the run, replacing the ice packs with new ones from
the Biokont-T container. Tomorrow is closeout day for this
experiment.]
FE-1 Nikolai Budarin transferred previous measurements of the
Molniya-SM experiment to HDD (hard disk drive) for return to Earth.
Using French LSO equipment, the experiment had its last run, from
Service Module (SM) window #3, on 4/10 and was dismantled on 4/14,
hopefully having caught some "sprites". [Objective of LSO was to
study rare optical phenomena occurring in the upper layers of Earth's
atmosphere, so-called "sprites" (i.e., puzzling glow phenomena
observed above thunderstorm clouds). LSO was originally part of
Claudie Haigneré's French "Andromeda" payload package of taxi mission
3S that could not be performed as planned during Increment 4 due to
an ISS flight attitude conflict. LSO research objectives are very
similar to those of the Russian Molniya-SM, being about storm
phenomena and other related events in the Earth's equatorial
regions. LSO was controlled from the French EGE-1 laptop, which
needs to be loaded with orbital sighting predictions using an
up-to-date NORAD tracking TLE (two-line element).]
Budarin deinstalled and removed the LKT local temperature sensor
commutator (TA251M1B) of the BITS onboard telemetry system and its
ROM (read-only memory) unit (TA765B) from the Soyuz TMA-1/5S Orbital
Module (OM), now no longer required (OM to be jettisoned tomorrow
evening) and to be reused in the future.
As part of the CDRA (carbon dioxide removal assembly) handover
activities, CDR Ken Bowersox and FE-1-7 Edward Lu conducted a
procedure review. replaced one of the ASV (air supply valves) filters
and tightened the CDRA pump QD (quick disconnect) coupling. More
work will be performed some time after undock. [After running
nominally, CDRA yesterday came up failed during a half-cycle (
switchover between channels). Data analysis isolated ASV #1 as the
"culprit" this time. When the valves were enabled, ASV-1 came up
"indeterminate", which has not been seen before (probably due to the
fact that the present high-rate data do not compare well to the
previous low-rate data). After MCC-H/ECLSS commanded a BIT (built-in
test), the valve recovered its functionality, and CDRA resumed
nominal ops. Data analysis continues to see whether this could be a
prelude to a valve failure.]
After double-checking the connectivity of the UOP/DCP (utility outlet
panel/display & control panel) power bypass cable for the
robotics/MSS (mobile service system), Bowersox and Lu took the
Canadian SSRMS (space station remote manipulator system) on its
scheduled familiarization run. [Robotics activities included initial
DCP switch tests, MSS camera settings, a joint OCAS (operator
commanded auto sequence) to a Lab PDGF (power & data grapple fixture)
pre-grapple position, followed by grapple, release and a single joint
maneuver to final stow position. SSRMS operations are not currently
planned for Expedition 7, but one never knows...]
Yesterday's resizing of the EMUs (extravehicular mobility units) for
Malenchenko and Lu went well. A don/doff demo session for the two
station residents is planned for later this month.
Nikolai Budarin conducted his daily checkup of the BIO-5
"Rasteniya-2/Lada-2" zero-G plant growth experiment. He copied
accumulated data to a memory card, performed video recording of the
plants (acacia-leaf peas) and cut three plants of medium size ("with
white flowers") for return to Earth, along with the video and PCMCIA
(portable computer memory card international adapter) card.
Budarin also conducted the atmospheric microbial air sampling
analysis of the Russian MO-21 "Ecosphera" air sampler and incubation
equipment, which he had set up yesterday. [MO-21 determines
sanitary/epidemiological conditions of the ISS atmosphere,
specifically the total bacterial and fungal microflora counts and
microflora composition according to morphologic criteria of
microorganism colonies.]
Later, Nikolai collected samples for analysis in the biosample kit of
the Russian BTKh-11 Biodegradatsiya ("biodegradation") experiment,
brought up on 5S, photographing the sampling ops with the Nikon F5
camera.
Budarin closed out the long-running Russian "Biorisk-MSV" experiment
and transferred the kit from the SM to Soyuz TMA-1 for return to
Earth. [Biorisk units, which originally arrived with Progress 9P in
late September '02, have been monitoring the cabin air.]
Malenchenko conducted his first regular (monthly) checkup of the
DC-1 "Pirs" docking compartment's circuit breakers (AZS) on the BVP
amp switch panel -- they should all be On -- and the 14 LEDs of the
fuses in fuse panels BPP-30 and BPP-36 (last time done: 4/18).
Yuri also checked out the IP-1 airflow sensors which are installed in
the various RS (Russian segment) hatchways (last time done: 4/16).
Nikolai Budarin performed another session of the Russian Uragan earth
observations program, using the Nikon D1 digital camera from SM
windows 5-8. The images, stored on an HDD, will be returned to Earth
on Soyuz 5S. [Targets of interest were the city of London and the
southern bank of the Thames River, the Rhine River and the largest
conurbation in the world (seven cities on the Rhine River/Ruhr Basin
merged into a single megalopolis 50ő30 km), open pit mines south of
Leipzig/Germany, the city of Prague, the Sudeten-Tatra-Carpathian
Mountains, the city of Uzhgorod, a detailed panorama of the Crimean
South Shore, low-oblique imagery along the Caucasian coastline, the
Borjomi-Bakuriani region on the Kura River (city of Bakuriani), and
low-oblique imagery of the Armenian territory.]
FE-1/SO Don Pettit performed the periodic transfer of accumulated
data files from the CMS HRM (crew medical systems/heart rate monitor)
to the MEC (medical equipment computer), then deleted them on the
HRM. Subsequently, he also transferred data files from the physical
exercise equipment (TVIS and RED) to the MEC.
Malenchenko completed the daily routine task of SOZh life support
systems maintenance and also prepared the daily IMS (inventory
management system) "delta" file for updating the IMS database, while
Ed Lu conducted the regular status checkup of autonomous Lab
payloads.
- Continued -
@Message posted automagically by IMTHINGS POST 1.30
---
* Origin: SpaceBase(tm) Pt 1 -14.4- Van BC Canada 604-473-9358 (1:153/719.1)SEEN-BY: 633/267 270 @PATH: 153/719 715 7715 140/1 106/2000 633/267 |
|
| SOURCE: echomail via fidonet.ozzmosis.com | |
Email questions or comments to sysop@ipingthereforeiam.com
All parts of this website painstakingly hand-crafted in the U.S.A.!
IPTIA BBS/MUD/Terminal/Game Server List, © 2025 IPTIA Consulting™.